two positive charges are located at a distance 0.58m apart on the a-axis. At the origin, the charge is 5.5x10^-8C, and at x=0.58m the charge is 3.3x10^-8C.At what point is the electric field equal to zero.
Answer #1

Let us consider that point is in between the charge particle somewhere at a distance d right to the charge placed at origin,

Now at that point electric field due to q1 and q2 must be equal.

k q1 / d^2 = k q2 / ( r - d) ^2

r/ d - 1 = sqrt(q2/q1)

0.58 / d - 1 = sqrt ( 3.3 / 5.5)

d = 0.3268 m

=======

Comment in case any doubt, will reply for sure.. Goodluck